OpenStudy (anonymous):

A cylindrical cardboard mailing tube has a diameter of 90 mm. and a height of 600 mm. A total of 20% of the tube’s capacity is filled with spherical packing foam. Each foam sphere has a radius of 15 mm. Approximately how many pieces of foam are in the tube? Use 3.14 to approximate pi and express your answer as a whole number.

OpenStudy (welshfella):

First find the volume of the cylinder :- 3.14*r^2 * h where h = height and r = radius of cylinder (note radius = 1/2 diameter) then find 20% of this volume and call it V2 next find the volume of each spherical piece of foam which is (4/3)* 3.14* r^3 where r= radius (15mm) finally divide this last volume into V2 to give you the number of spheres